You are given the following information about a closed economy with no government: Consumption = 115 + 0.6Y Investment = 550 Use the above information to answer the questions that follow:
Q.4.4 Is the equilibrium level of income also the full employment level of income? Explain your answer
1
Expert's answer
2020-09-22T15:27:09-0400

The equilibrium level of national income is a national income determined when aggregate expenditure equals national income. At this point injections in the circular flow equals withdrawals. From the given data, the equilibrium national income is calculated as: "Y = C + I"

"=> Y = 115 + 0.6Y + 550"

"=> Y - 0.6Y = 115 + 550"

"=> 0.4Y = 665"

"=> Y = \dfrac {665}{0.4}"


"=> Y = \$1,662.50"


Therefore, $1,662 is the equilibrium national income as its equates aggregate expenditure (C + I) to national income.


The equilibrium national income is not necessarily the full employment equilibrium. It can be higher or lower than the full employment output. If the equilibrium exists in the short or below the economy's production potential then it is lower than full employment. A lower than full employment equilibrium implies that the economy is producing below its production possibility curve; there are unemployed resources in the economy and the equilibrium do not represent potential output. Also, the equilibrium can be above full employment where the actual national output exceeds potential output. At such as equilibrium, prices are unstable and inflation exists.


Therefore, equilibrium national income do not necessarily represent full employment national output.